# Democratizing Large Language Models via Personalized Parameter-Efficient Fine-tuning

### v1

@@ Abstract.
Personalization in large language models (LLMs) is increasingly important, aiming to align LLM’s interactions, content, and recommendations with individual user preferences. Recent advances in LLM personalization have spotlighted effective prompt design, by enriching user queries with non-parametric knowledge through behavior history retrieval and textual profiles. However, these approaches were limited due to a lack of model ownership, resulting in constrained customization and privacy issues. Moreover, they often failed to accurately capture user behavior patterns, especially in cases where user data were complex and dynamic. To address these shortcomings, we introduce One PEFT Per User (OPPU) [^1] The code is available at https://github.com/TamSiuhin/OPPU, which employs personalized parameter-efficient fine-tuning (PEFT) modules, to store user-specific behavior patterns and preferences. By plugging in users’ personal PEFT parameters, they can own and use their LLMs personally. OPPU integrates parametric user knowledge in the personal PEFT parameters with the non-parametric knowledge acquired through retrieval and profile. This integration adapts individual LLMs to user behavior shifts. Experimental results demonstrate that OPPU significantly outperforms existing prompt-based methods across seven diverse tasks in the LaMP benchmark. Further in-depth studies reveal OPPU’s enhanced capabilities in handling user behavior shifts, modeling users at different active levels, maintaining robustness across various user history formats, and displaying versatility with different PEFT methods.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Personalization refers to mining users’ behavior history, and therefore tailoring and customizing a system’s interactions, content, or recommendations to meet specific needs, preferences, and characteristics of individual users (Tan and Jiang, 2023; Chen et al., 2023; Chen, 2023). By adapting to each user’s preferences, personalization systems enhance user experience, increasingly getting vital in areas like content recommendation (Qian et al., 2013; Li et al., 2023a; Wu et al., 2023; Baek et al., 2023), user simulation (Dejescu et al., 2023; Zhang and Balog, 2020), personalized chatbots (Srivastava et al., 2020; Ma et al., 2021; Qian et al., 2013), user profiling (Gu et al., 2020; Gao et al., 2023), healthcare (Johnson et al., 2021; Goldenberg et al., 2021), and education (Alamri et al., 2021; Pratama et al., 2023).
Large language models (LLMs) display emergent abilities not seen in smaller models (Wei et al., 2022a; Lu et al., 2023), as they have billions of parameters and are trained on vast corpora. These abilities include step-by-step reasoning (Wei et al., 2022b), in-context learning (Min et al., 2022), and instruction following (Wei et al., 2021). However, existing LLMs predominantly follow the “one-size-fits-all” paradigm. They are generally trained on extensive, domain-agnostic datasets, which limits their effectiveness in meeting the specific needs and preferences of individual users (Chen et al., 2023; Salemi et al., 2023). Therefore, the challenge of integrating the strong generative capabilities of LLMs with the tailored requirements of individual users has emerged as a significant area of research (Jang et al., 2023; Li et al., 2023b).
Existing works on personalizing LLMs have predominantly concentrated on the development of prompt templates. These prompt-based personalization methods fall into three categories: vanilla personalized prompt, retrieval-augmented personalized prompt, and profile-augmented prompt. The vanilla personalized prompt approach leverages the in-context learning capability of LLMs, utilizing the user’s entire or random sampled behavior history as contextual examples (Zhiyuli et al., 2023; Dai et al., 2023; Christakopoulou et al., 2023). Considering the growing length of user behavior history and the limited context window of LLMs, some studies have applied retrieval methods to select the most relevant part of user behavior history to enhance LLM personalization (Salemi et al., 2023; Mysore et al., 2023). Besides the retrieval, some techniques explicitly generate user preferences and profiles in natural language to augment LLMs’ input. For instance, Richardson et al. (2023) proposed to employ instruction-tuned LLMs, e.g., ChatGPT, to summarize user preferences and behavior patterns based on their history content.
@@
Figure 1. LLM ownership and behavior shift are two challenges that developing personalized LLMs has to face. Ownership emphasizes that the model needs to be owned by individual user to enhance customization and privacy. Behavior shift adaption refers to the LLMs’ ability to effectively generalize and adapt to emerging new patterns in user behaviors.
@@
Despite much research progress has been made in LLM personalization, existing methods face ownership and behavior shift challenges, which are illustrated in Figure 1:
@@
  • Ownership: Existing methods are processed in a centralized way, where user history is encoded in a personalized prompt and processed by centralized LLMs. This paradigm limits the model’s customization and ability to provide deep, personalized experiences tailored to individual users. Moreover, when using a centralized model, users often have to share personal data with the service provider, which raises concerns about how user data are stored, used, and protected.
@@
  • Behavior Pattern Generalization: As is revealed by Shi et al. (2023), LLMs can be easily distracted by irrelevant context information that retrieval can hardly avoid. In the realm of LLM personalization, where the retrieval corpus was confined to a specific user’s behaviors, retrieval augmentation might underperform, especially when the user’s past behaviors did not closely mirror the patterns needed for the query at hand.
@@
In light of these challenges, we propose One PEFT Per User (OPPU), where each user is equipped with a personalized, parameter-efficient fine-tuning (PEFT) module. Characterized by PEFT’s plug-and-play functionality and the minimal weight of updated parameters (typically less than 1% of the base LLM), OPPU facilitates LLM ownership and exhibits superior generalization in scenarios of user behavior shifts. By fine-tuning the PEFT module with the user’s personal behavior history, the personalized PEFT parameters encapsulate behavior patterns and preferences. This process, when integrated into base LLMs, allows users to obtain their private LLMs, ensuring LLM ownership and enhancing model customization. Furthermore, as is revealed by Gupta et al. (2024), fine-tuning LLMs is more effective than retrieval augmentation when the retrieved instances are not highly relevant to the query. The fine-tuned personal LLMs in OPPU are adept at capturing complex behavior patterns and thus capable of understanding new behaviors with less reliance on highly relevant history data.
Experimental results show that OPPU achieves state-of-the-art performance on all seven public tasks in the Language Model Personalization (LaMP) benchmark (Salemi et al., 2023). Additional studies highlight the importance of integrating non-parametric user knowledge, sourced from retrieved user history, with parametric user knowledge from personal PEFT parameters. Notably, in scenarios of user behavior shift, where the user history is less relevant to the current user query at hand, OPPU significantly outperforms retrieval-based methods. Moreover, OPPU exhibits strong resilience against varying user history formats and demonstrates versatility across different PEFT methods, among other advantages.
@@
To summarize, the contribution of OPPU lies in its pioneering approach to PEFT-based LLM personalization. Each user (or user cohort) benefits from a personal PEFT module, which not only ensures LLM ownership but also significantly improves the model’s ability to adapt to shifts in user behavior. The superiority of OPPU is evidenced by state-of-the-art performance across seven tasks in the LaMP benchmark. By introducing this innovative parametric-based personalization technique, OPPU opens up new opportunities in the realm of democratizing personalized LLMs.
2. Related Work
2.1. Personalization of LLMs
The thrust of existing LLM personalization research is centered on designing prompt that incorporate historical user-generated content and behavior. These approaches help LLMs understand users’ preferences, tailoring responses to individual needs (Tan and Jiang, 2023; Chen et al., 2023). The endeavors towards personalized LLMs mainly fall into three categories: vanilla personalized prompts, retrieval-augmented personalized prompts, and profile-augmented personalized prompts.
In the vanilla personalized prompt category, researchers use in-context and few-shot learning to encode either complete or a sample of user behavior history as contextual examples. For instance, Dai et al. (2023) and Kang et al. (2023) encode the user’s personal rating history as few-shot demonstration examples. Liu et al. (2023b) supply LLMs with user’s interaction history related to specific tasks to help LLMs generate personalized content. PerSE (Wang et al., 2023) design prompt to do personalized story assessment by presenting a few exemplary reviews from the review. BookGPT (Zhiyuli et al., 2023) employs a few-shot prompt strategy to make LLMs understand the correlation between book content and personalized prompting. Moreover, some research works (Christakopoulou et al., 2023; Zhiyuli et al., 2023) also discovered a long user history would bring better performance. To manage the growing user behavior data and LLMs’ limited context windows, the retrieval-augmented personalized prompt approach has emerged. For instance, LaMP (Salemi et al., 2023) introduces a retrieval-augmented method to obtain the most relevant content in the user’s behavioral history and incorporate it into the prompt. AuthorPred (Li et al., 2023b) utilizes retrieve relevant past user-written documents for personalized text generation. Pearl (Mysore et al., 2023) proposes a generation-calibrated retriever to select historic user-authored documents for prompt augmentation. Moving beyond simple retrieval, some researchers summarize user preferences and behavior patterns into natural language profiles for input query augmentation, termed profile-augmented personalized prompts. Richardson et al. (2023) use the instruction-tuned LLMs to generate an abstract summary of user history data, augmenting retrieval-based personalization methods. ONCE (Liu et al., 2023a) employs LLMs to generate users’ topics and regions of interest based on their browsing history as user profiles, aiding LLMs in preference capture in downstream tasks. There is also another line of work focusing on designing personalized alignment methods via parameter merging (Jang et al., 2023) and personalized reward model (Cheng et al., 2023).
Previous works largely hinge on prompt design, integrating retrieval and user profiles. However, existing approaches are limited by model ownership and users’ behavior shifts. Our OPPU introduces personalization at the parametric level, via the personal PEFT module, which ensures model ownership and superior generalization, especially in cases with less relevant user history for retrieval.
2.2. Parameter-Efficient Fine-tuning (PEFT)
With the exponentially growing parameters in LLMs, fine-tuning all parameters is expensive (Liu et al., 2022a; Xu et al., 2023; Gupta et al., 2024). To mitigate this gap, a few lightweight alternatives, known as parameter-efficient fine-tuning (PEFT) have been proposed to update only a small number of extra parameters while keeping the pretrained weights frozen to save the computes (He et al., 2021; Fu et al., 2023). For example, adapter tuning (Houlsby et al., 2019) injects learnable parameters in the models’ each feedforward layer, and only the plug-in parameters are updated at the fine-tuning stage. Inspired by the success of discrete textual prompt (Sanh et al., 2022; Wang et al., 2022), prefix tuning (Li and Liang, 2021) and prompt tuning (Lester et al., 2021) are proposed to learn the optimized prompt and prefix for specific use via fine-tuning. LoRA (Hu et al., 2021) proposes to add low-rank matrices on pretrained weights to approximate parameter updates. $\rm{(IA)^{3}}$ (Liu et al., 2022b) plugs in learned vectors to scale the activation in the attention mechanism for efficient fine-tuning. These approaches achieve comparable performance to full parameter fine-tuning by updating less than 1% of the original LLM parameters. Besides parameter savings, PEFT methods are effective in combating catastrophic forgetting (Pfeiffer et al., 2021) and robust to out-of-distribution samples (Li and Liang, 2021).
The small number of updated parameters and plug-and-play nature of PEFT make it an ideal solution for efficient LLM personalization and constitute model ownership. Our work pioneers the concept of storing user history within personal PEFT parameters. Each user is equipped with a unique, easily integrable PEFT module, serving as a key to their own personalized LLMs.
@@ 2.2. Parameter-Efficient Fine-tuning (PEFT)
Figure 2. Overview of our proposed OPPU, where each user is equipped with a personal PEFT module and plug-in base LLMs to get their individual LLM. Beyond parametric personalization via PEFT, OPPU is also compatible with the non-parametric user knowledge via retrieval and profile augmentation.
3. Proposed Approach
3.1. Research Problem Formulation
Generative language models generally take an input token sequence $x$ and output a sequence of tokens $y$ that follows the $x$. For the personalization of LLMs, happening at time $t$, the language model’s output $y_{u}$ specifically for user $u$ is conditioned on both input $x_{u}$ and the user $u$’s behavior history $\mathcal{H}_{u}$. To elaborate, $\mathcal{H}_{u}^{t}=\{h_{u}^{t_{i}}\},t_{i}<t$, where the user’s history data $\mathcal{H}_{u}$ contains all user behavior $h^{t_{i}}$ happened before query time $t$. More specifically, the user behavior $h_{u}^{t_{i}}$ may consist of $(x_{u}^{t_{i}},y_{u}^{t_{i}})$ pairs, mirroring the task-specific query-answer format $(x_{u},y_{u})$. Alternatively, $h_{u}^{t_{i}}$ could be text sequences that provide context for the user’s behavior patterns without necessarily conforming to the task’s format. Overall, each data entry contains three components: input sequence $x_{u}$ given by the user, user’s history behavior $\mathcal{H}_{u}$ that contains the user’s behavior that happened before the input query, and a target output $y_{u}$ specifically for user $u$ that the model is expected to produce.
3.2. Base LLMs Task Adaption
Given that off-the-shelf LLMs are not inherently equipped for personalization tasks, we align with the methods of LaMP (Salemi et al., 2023) and Richardson et al. (2023) by fine-tuning LLMs for fair comparison. In this section, we will describe the process of developing base LLMs, a critical step to enhance their general capability in comprehending and executing the specific personalization tasks required, without specific user’s preference or bias.
Non-personalized base LLM ($\mathrm{LLM_{NP}}$) ignores the user’s behavior history and serves as a baseline. It is fine-tuned on $\{(\phi_{p}(x_{u}),y_{u})\}$ data pairs, where $\phi_{p}$ is the prompt construction function. This approach only contains the task-related data and ignores the user behavior history, therefore making it non-personalized.
Retrieval-augmented base LLM ($\mathrm{LLM_{RAG}}$) is fine-tuned on input that includes the retrieved top-$k$ relevant user history items, which the input sequence $x^{\prime}$ can be defined as:
(1) $\overline{x_{u}}=\phi_{p}(x_{u},\mathcal{R}(x_{u},\mathcal{H}_{u},k)),$
where $\mathcal{R}$ denotes the retriever, $\phi_{p}$ denotes the prompt construction function. This approach aims to enhance personalization by integrating the most pertinent user history items directly into the LLM’s prompt.
Profile-Augmented Base LLM ($\mathrm{LLM_{PAG}}$) enhances its input by appending a natural language user profile that describes the user’s preferences and behavior patterns. This profile can be added to the non-personalized and retrieval-augmented prompts. The input $x^{\prime}$ to LLM could be denoted as:
(2) $\overline{x_{u}}=\phi_{p}(x_{u},\mathcal{R}(x_{u},\mathcal{H}_{u},k),s_{u}),$
where $s_{u}$ is a textual user profile that describes user’s preference and behavior patterns that are generated by an instruction-tuned LLM (Vicuna (Chiang et al., 2023) or ChatGPT (Ouyang et al., 2022)) based on the user history data.
To make this process more computationally efficient, we adopt the low-rank adaptation (LoRA) (Hu et al., 2021) for base modeling task adaption that only updates about 0.5% of external parameters compared to the total LLM parameter size. After training the LoRA, we merge the LoRA parameters to the base model to get base LLMs equipped with the capabilities for the corresponding tasks.
3.3. One PEFT Per User (OPPU)
After obtaining the base model that comprehends the responding task, in real-world deployment, users can only assess the parameter of the base model and their personal behavior history data, controlling the privacy risks. In this section, we will introduce how to make personalized LLMs via parameter-efficient fine-tuning (PEFT), and its integration with non-parametric personalization, i.e., retrieval-augmentation and profile-augmentation. To better understand the parametric personal knowledge in PEFT parameters and non-parametric personalized knowledge in the prompt, we decompose OPPU into parametric knowledge only and the integration of both parametric and non-parametric personalized knowledge.
3.3.1. Inject Parametric Personalized Knowledge via PEFT
We first explore injecting the parametric personalized knowledge into base LLMs and exclude the non-parametric augmentation method to focus purely on the impact of parametric modifications. In this scenario, the user history data are all stored in the PEFT module’s parameters thus enhancing privacy preservation. For each user $u$, OPPU would maintain a private PEFT module $\mathrm{PEFT}{u}$, tailored to capture and adapt to the user’s behavioral patterns as reflected in their historical data. More specifically, consider a user $u$ with history $\mathcal{H}_{u}$, the personalized PEFT module is plugged in base LLM ${LLM_{NP}}$ and optimized on the user’s personal history data. The input sequence $x^{\prime}$ for LLMs is $x_{u}^{\prime}=\phi_{p}(x^{t_{i}}_{u})$ and optimized using the standard cross entropy loss against the corresponding output $y^{t_{i}}_{u}$, where $(x^{t_{i}}_{u},y^{t_{i}}_{u})\in\mathcal{H}_{u}$. This procedure could capture the user behavior pattern presented in all user behavior history more comprehensively store it in plug-and-play personal PEFT parameters, constituting a personalized LLM owned by users.
3.3.2. Integrating Non-Parametric & Parametric Knowledge
Inspired by recent research on debating between fine-tuning versus retrieval-augmented generation (Gupta et al., 2024), as well as insights into the integration of non-parametric knowledge through retrieval and parametric knowledge via PEFT, we proceed to investigate the integration of our personalized PEFT parametric user knowledge with non-parametric user knowledge from both retrieval and profile augmentation methods.
Integrating with Retrieval Augmentation
Our first step involves integrating the personal PEFT module with retrieval augmentation, in which the $\mathrm{LLM_{RAG}}$ acts as the base LLM and each user’s plug-in PEFT parameters are updated on the following input sequence $x^{\prime}$:
(3) ${{x}_{u}^{t_{i}}}^{\prime}=\phi_{p}(x_{u}^{t_{i}},\mathcal{R}(x_{u}^{t_{i}},\mathcal{H}_{u}^{t_{i}},k)),$
where $\phi{p}$ denotes the prompt construction function and $\mathcal{R}$ represents the retriever. The retriever selects the top $k$ relevant history items from the user’s behavior history corpus $\mathcal{H}_{u}^{t_{i}}$ in response to the user history query $x_{u}$. It’s important to note that the retrieval is based on past behavior that occurred before the current query $x_{u}^{t_{i}}$. The PEFT parameters are updated using cross entropy loss against the corresponding historical user response $y_{u}^{t_{i}}$.
Integrating with User Profile Augmentation
Following Richardson et al. (2023), we also incorporate profile augmentation into our method for injecting non-parametric knowledge into prompts. In this scenario, the personal PEFT parameter $\mathrm{PEFT_{u}}$ is plugged in $\mathrm{LLM_{PAG}}$ and optimized using the input ${{x}_{u}^{t_{i}}}^{\prime}$:
(4) ${{x}_{u}^{t_{i}}}^{\prime}=\phi_{p}(x_{u}^{t_{i}},\mathcal{R}(x_{u}^{t_{i}},\mathcal{H}_{u}^{t_{i}},k),s_{u}),$
which includes the user’s profile $s_{u}$ on the basis of retrieval augmentation. This profile $s_{u}=LLM(\mathcal{H}_{u})$ is a summary of the user’s preferences and behavior patterns, generated by an LLM that can follow human instruction.
By integrating the user knowledge in both private PEFT module and non-parametric retrieval and profile methods, OPPU is better equipped to understand a user’s behavior patterns. This integrated approach enhances the model’s ability to generalize these patterns to new and emerging user behaviors.
3.3.3. Adapting to Behavior History Beyond Task Conformity
It is worth mentioning that the user behavior history does not always align neatly with the format of a user’s query. For instance, in personalized tweet paraphrasing tasks, where to input a sequence of text $x_{u}$, the desired output is the corresponding paraphrased text $y_{u}$ concerning user history $\mathcal{H}_{u}$. However, $\mathcal{H}_{u}$ often comprises only the user’s historical tweets $\{y_{u}^{t_{i}}\}\in\mathcal{H}_{u}$, not the $(x_{u}^{t_{i}},y_{u}^{t_{i}})$ pairs. In scenarios where user history does not directly provide preferences in the specific task format, we find tuning a personal PEFT could also provide LLMs with more context and benefit the personalization task performance. Specifically, we update the personal PEFT parameters using unsupervised pretraining loss to predict next tokens over the user history sequences $h_{u}^{t_{i}}\in\mathcal{H}_{u}$. This approach infuses personal historical knowledge into the LLM via the integrated PEFT parameters.
Overall, we envision the proposed OPPU as a versatile LLM personalization framework, where each user possesses their own PEFT parameters that contain their personal behavior history and preference. By plugging their personal PEFT parameters into the base LLMs, users could get their personalized LLMs, while achieving a better understanding of users’ personality from the parametric dimension.
4. Experiments
4.1. Experimental Settings

### v2

@@ Abstract
Personalization in large language models (LLMs) is increasingly important, aiming to align the LLMs’ interactions, content, and recommendations with individual user preferences. Recent advances have highlighted effective prompt design by enriching user queries with non-parametric knowledge through behavior history retrieval and textual profiles. However, these methods faced limitations due to a lack of model ownership, resulting in constrained customization and privacy issues, and often failed to capture complex, dynamic user behavior patterns. To address these shortcomings, we introduce One PEFT Per User (OPPU) [^1] The code is available at https://github.com/TamSiuhin/OPPU, employing personalized parameter-efficient fine-tuning (PEFT) modules to store user-specific behavior patterns and preferences. By plugging in personal PEFT parameters, users can own and use their LLMs individually. OPPU integrates parametric user knowledge in the personal PEFT parameters with non-parametric knowledge from retrieval and profiles, adapting LLMs to user behavior shifts. Experimental results demonstrate that OPPU significantly outperforms existing prompt-based methods across seven diverse tasks in the LaMP benchmark. Further studies reveal OPPU’s enhanced capabilities in handling user behavior shifts, modeling users at different activity levels, maintaining robustness across various user history formats, and displaying versatility with different PEFT methods.
Democratizing Large Language Models via
Personalized Parameter-Efficient Fine-tuning
Zhaoxuan Tan, Qingkai Zeng, Yijun Tian, Zheyuan Liu, Bing Yin, Meng Jiang University of Notre Dame, Amazon.com Inc. {ztan3, mjiang2}@nd.edu
1 Introduction
@@
Personalization refers to mining users’ behavior history, and therefore tailoring and customizing a system’s interactions, content, or recommendations to meet specific needs, preferences, and characteristics of individual users Tan and Jiang (2023); Chen (2023). By adapting to each user’s preferences, personalization systems enhance user experience, increasingly getting vital in areas like content recommendation Qian et al. (2013); Wu et al. (2023); Baek et al. (2023), user simulation Dejescu et al. (2023), personalized chatbots Srivastava et al. (2020); Ma et al. (2021), user profiling Gu et al. (2020); Gao et al. (2023), healthcare Goldenberg et al. (2021), and education Pratama et al. (2023).
@@
Figure 1: LLM ownership and behavior shift are two challenges that developing personalized LLMs has to face. Ownership emphasizes that the model needs to be owned by individual user to enhance customization and privacy. Behavior shift adaption refers to the LLMs’ ability to effectively generalize and adapt to emerging new patterns in user behaviors.
Large language models (LLMs) display emergent abilities not seen in smaller models Wei et al. (2022); Lu et al. (2023), as they have billions of parameters and are trained on vast corpora. However, existing LLMs predominantly follow the “one-size-fits-all” paradigm. They are generally trained on extensive, domain-agnostic datasets, which limits their effectiveness in meeting the specific needs and preferences of individual users Chen et al. (2023). Therefore, the challenge of integrating the strong generative capabilities of LLMs with the tailored requirements of individual users has emerged as a significant area of research Li et al. (2023).
Existing works on personalizing LLMs have predominantly concentrated on developing prompt templates, which fall into three categories: vanilla, retrieval-augmented, and profile-augmented personalized prompts. The vanilla personalized prompt approach leverages the in-context learning capability of LLMs, utilizing the user’s entire or randomly sampled history as contextual examples Dai et al. (2023); Zhiyuli et al. (2023). Considering the growing length of user behavior history and the limited LLM context length, some studies applied retrieval methods to select the most relevant part of user behavior history to enhance LLM personalization Mysore et al. (2023). Besides the retrieval, some techniques explicitly generate user preferences and profiles in natural language to augment LLMs’ input Richardson et al. (2023).
@@
Despite much research progress has been made in LLM personalization, existing methods face ownership and behavior shift challenges (Fig. 1):
@@
  • Ownership: Existing methods are processed centralized, where user history is encoded in a personalized prompt and processed by centralized LLMs. This paradigm limits the model’s customization and ability to provide deep, personalized experiences tailored to individual users. Moreover, when using a centralized model, users often have to share personal data with the service provider, which raises concerns about how user data are stored, used, and protected.
@@
  • Behavior Pattern Generalization: As is revealed by Shi et al. (2023), LLMs can be easily distracted by irrelevant context information that retrieval can hardly avoid. In LLM personalization, where the retrieval corpus is confined to a specific user’s behaviors, retrieval augmentation might underperform, especially when the user’s past behaviors do not closely mirror the patterns needed for the query at hand.
@@
In light of these challenges, we propose One PEFT Per User (OPPU), equipping each user with a personalized, parameter-efficient fine-tuning (PEFT) module. Characterized by PEFT’s plug-and-play functionality and the minimal weight of updated parameters (typically less than 1% of the base LLM), OPPU facilitates LLM ownership and enhances generalization in scenarios of user behavior shifts. By fine-tuning the PEFT module with the user’s personal behavior history, the personalized PEFT parameters encapsulate behavior patterns and preferences. This process, when integrated into base LLMs, allows users to obtain their private LLMs, ensuring LLM ownership and enhancing model customization. Furthermore, as is revealed by Gupta et al. (2024), fine-tuning LLMs is more effective than retrieval augmentation when the retrieved instances are not highly relevant to the query. The fine-tuned personal LLMs in OPPU are adept at capturing complex behavior patterns and thus capable of understanding new behaviors with less reliance on highly relevant history data. Experimental results show that OPPU outperforms all baselines on seven public tasks in the Language Model Personalization (LaMP) benchmark Salemi et al. (2023). Additional studies emphasize the importance of integrating non-parametric user knowledge from retrieved history with parametric knowledge from personal PEFT parameters. In scenarios of user behavior shifts, where history is less relevant, OPPU significantly outperforms retrieval-based methods. Moreover, OPPU is resilient to varying user history formats and demonstrates versatility across different PEFT methods, among other advantages.
@@
To summarize, the contribution of OPPU lies in its pioneering approach to PEFT-based LLM personalization. Each user (or user cohort) benefits from a personal PEFT module, which not only ensures LLM ownership but also significantly improves the model’s ability to adapt to shifts in user behavior. The superiority of OPPU is evidenced by state-of-the-art performance across seven tasks in the LaMP benchmark. By introducing this innovative parametric-based personalization technique, OPPU opens up new opportunities in democratizing personalized LLMs.
@@
Figure 2: Overview of our proposed OPPU, where each user is equipped with a personal PEFT module and plug-in base LLMs to get their individual LLM. Beyond parametric personalization via PEFT, OPPU is also compatible with the non-parametric user knowledge via retrieval and profile augmentation.
2 Preliminaries
2.1 Research Problem Formulation
For personalizing LLMs at time $t$, the output $r_{u}$ for user $u$ is conditioned on both input $q_{u}$ and the user’s behavior history $\mathcal{H}_{u}$. Specifically, $\mathcal{H}_{u}=\{h_{u}\}$, includes all user behaviors $h_{u}$ before time $t$. User behavior $h_{u}$ may consist of $(x_{u},y_{u})$ pairs, aligning with the task-specific query-answer format $(q_{u},r_{u})$, or plain text sequences $x_{u}$ providing context for behavior patterns. We aim to obtain personalized parameters $\Theta_{u}$ for each user $u$.
2.2 Base LLMs Task Adaption
Given that off-the-shelf LLMs are not inherently equipped for personalization tasks, we follow the methods of LaMP Salemi et al. (2023) by fine-tuning LLMs for fair comparison. This section outlines the development of base LLMs with a set of held-out users to enhance their general capabilities for personalization tasks without involving target user preferences. Specifically, we provide three alternatives: base LLM $\Theta^{\mathrm{(B)}}$ that only involves task related data, retrieval-augmented base LLM $\Theta^{\mathrm{(R)}}$ that augment input with top-$k$ relevant user history, and profile-augmented base LLM $\Theta^{\mathrm{(P)}}$ that involves textual user profiles as input. Note that introducing RAG and PAG means users would expose their historical data or profiles to a centralized LLM, potentially affect the model ownership. For users prioritizing privacy and ownership, OPPU without retrieval avoids revealing user data to service providers. Conversely, those seeking optimal performance and consent to reveal data to centralized LLMs should opt for RAG or PAG. The fine-tuning objectives of three base models are: